Circular No. 7668

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

PROBABLE SUPERNOVA IN ANONYMOUS GALAXY P. Challis, Harvard-Smithsonian Center for Astrophysics, on behalf of

the High-Z Supernova Search Team (cf. IAUC 6160, 6646), using methods inspired by the SNAZ (SuperNovae at High Z) Team at the Space Telescope Science Institute, reports the discovery of a probable supernova on WFPC2 images taken with the Hubble Space Telescope. The apparent supernova was found by differencing images taken on June 20 UT with earlier WFPC2 data taken over six epochs during Nov.–Dec. 2000. The new object appears in five filters during the 3-orbit (∼ 5-hr) observation. It does not exhibit the parallax motion of a solar-system object. The supernova is offset 0′′.4 southeast of the center of an elongated faint (F814W magnitude ∼ 24.1) galaxy. The position of the new object is α = 2^{h}30^{m}17^{s}.62, δ = +1^{o}03′56′′.4 (equinox 2000.0; uncertainty ± 0′′.04). A finder chart is available at http://cfa-

www.harvard.edu/cfa/oir/Research/supernova/HighZ.html. Photom-

etry from the WFPC2 data yields the following supernova magnitudes:

[F450W] = 26.00 ± 0.30, [F555W] = 25.07 ± 0.16, [F675W] = 23.91 ± 0.06, [F814W] = 23.59 ± 0.06, [F850LP] = 23.45 ± 0.16. Further observations with HST will be coordinated by the SNAZ Team. Groundbased observations should be possible in the R, I, and Z bands.

V2274 CYGNI = NOVA CYGNI 2001 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V2274 Cyg has been given to this nova (cf. IAUC 7666).

T. Kato, Kyoto University, reports the following prediscovery limiting magnitudes (nova not visible): May 12.775 UT, [14.0 (A. Takao, Kitakyushu, Japan; unfiltered CCD); 19.665, [11.9 (K. Kanatsu, Matsue, Shimane, Japan; T-Max 400 film); June 15.621, [11.9 (Kanatsu); July 12.613, [11.6 (Kanatsu).

E. O. Waagen, AAVSO, reports the following additional CCD V magnitudes in the AAVSO International Database: July 19.072 UT, 12.2 (A. Price, Waltham, MA); 19.090, 11.95 (Walker); 19.138, 11.87 (Walker); 19.159, 11.93 (Walker); 19.292, 11.86 (S. O’Connor, Montreal, QC); 20.058, 11.88 (Walker); 20.086, 11.81 (Walker); 20.096, 11.87 (Walker); 20.102, 12.2 (Price); 20.262, 12.0 (R. Royer, Lakewood, CA).

2001 July 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green